melbohme:

When solving a system of equations, Jared found y = x + 10 for one equation and substituted x + 10 for y in the other equation. Nicole found x = y – 10 for the same equation and substituted y – 10 for x in the other equation. Who is correct? Explain.

darkknight:

They both are essentially correct in the process. In substitution you solve for 1 variable and plug in that value into the other equation. For example if y=x+10, you can plug that in for y, Now y=x+10, we can subtract 10 on both sides and get x = y-10. Which can also be used to substitute for x in the other equation

melbohme:

thank you !!